# SAGE: Strategy-Adaptive Generation Engine for Query Rewriting

### v1

@@ 5.1 The Impact of Strategic Guidance
Table 3: Ablation study on the impact of strategic guidance. We compare the performance of a standard RL baseline (DeepRetrieval) against our SAGE framework, which incorporates expert-crafted strategies. The results highlight the significant gains in both effectiveness (NDCG@10) and efficiency (Avg. Tokens) brought by our strategic guidance.
@@ 5.3 The Critical Role of the Exploration Penalty
Conservative Prompt ("keep unchanged") 0.671 0.509
Proactive Exploration Prompt 0.694 0.951
Proactive Prompt + Penalty ($p=0.02$) 0.692 0.998
To validate the effectiveness of these mechanisms, we conducted an ablation study using the SAGE-SCS model on HotpotQA. The results, shown in Table 4, clearly demonstrate the value of our approach: simply replacing the conservative prompt with our proactive exploration prompt substantially increases both NDCG@10 and the modification rate, highlighting the critical role of encouraging exploration. However, the use of an exploration penalty reveals a critical trade-off. While it drives the modification rate to nearly perfect levels (0.998), we observe a slight decrease in NDCG@10. This suggests that in cases where the original query is already optimal, enforcing modifications via a penalty may slightly impair peak performance. Thus, although the penalty is highly effective for maximizing exploration and mitigating reward hacking, its application should be carefully balanced against the specific task’s need for exhaustive exploration versus optimal end performance.
@@ 6 Conclusion
In this work, we introduced SAGE, a novel framework designed to address the black-box nature and inefficient exploration of conventional RL-based query rewriting. By deeply integrating expert-crafted strategies and novel reward shaping mechanisms (SCS and CRS) into an RL loop, SAGE not only achieves a new state-of-the-art in retrieval effectiveness but also exhibits a compelling emergent efficiency, substantially reducing inference costs. Our findings establish that strategy-guided RL is a powerful framework for developing retrieval systems that are not only more effective and efficient, but also more transparent and controllable, opening promising avenues for future research in areas like automatic strategy discovery.
@@ A.1 Strategy-Guided Prompting
As referenced in Section 3.2, we present detailed strategies and the complete prompt used to guide our policy model in Figure 5.
@@ A.2 SGP Performance with Contriever
To confirm that our findings are robust and not specific to a single retriever, we replicated the experiment using Contriever Izacard et al. [2021]. As detailed in Fig 6, our strategy again demonstrates a significant performance margin over the methods proposed by DMQR-RAGLi et al. [2024].
Figure 6: Performance scaling laws for query rewriting using the Contriever Izacard et al. [2021] retriever across four benchmark datasets: HotpotQA Yang et al. [2018], FEVER Thorne et al. [2018], NFCorpus Boteva et al. [2016], and SciFact Wadden et al. [2020]. Each subplot compares the upper-bound performance (NDCG@10) of our Strategy-Guided Prompting (SGP) against the strategies from Li et al. [2024], evaluated via a Best-of-N methodology across the Qwen3 model series Yang et al. [2025b]. The results demonstrate the superior scaling potential and higher performance ceiling of our approach.
A.3 Detailed Training and Response Length Curves

### v2

@@ 3.2 Comparison with State-of-the-Art Baselines
Table 2: Ablation study on the impact of strategic guidance. We compare the performance of a standard RL baseline against our SAGE framework, which incorporates expert-crafted strategies. The results highlight the significant gains in both effectiveness (NDCG@10) and efficiency (Avg. Tokens) brought by our strategic guidance.
4 Ablation Study
4.1 The Impact of Strategic Guidance
To provide a more intuitive understanding, Fig.  4 presents a direct case comparison. The output from the model trained without strategies is generic and fails to capture the query’s nuance, whereas the output from SAGE is scientifically precise and far more effective for retrieval.
Conservative Prompt ("keep unchanged") 0.671 0.509
Proactive Exploration Prompt 0.694 0.951
Proactive Prompt + Penalty ($p=0.02$) 0.692 0.998
@@ 4.3 The Critical Role of the Exploration Penalty
To validate the effectiveness of these mechanisms, we conducted an ablation study using the SAGE-SCS model on HotpotQA. The results, shown in Table 3, clearly demonstrate the value of our approach: simply replacing the conservative prompt with our proactive exploration prompt substantially increases both NDCG@10 and the modification rate, highlighting the critical role of encouraging exploration. However, the use of an exploration penalty reveals a critical trade-off. While it drives the modification rate to 0.998, we observe a slight decrease in NDCG@10. This suggests that in cases where the original query is already optimal, enforcing modifications via a penalty may slightly impair peak performance. Thus, although the penalty is highly effective for maximizing exploration and mitigating reward hacking, its application should be carefully balanced against the specific task’s need for exhaustive exploration versus optimal end performance.
@@ 5 Conclusion
In this work, we introduced SAGE, a novel framework designed to address the black-box nature and inefficient exploration of conventional RL-based query rewriting. By integrating expert-crafted strategies and novel reward shaping mechanisms (SCS and CRS) into the RL loop, SAGE achieves state-of-the-art retrieval effectiveness and emergent efficiency, substantially reducing inference costs. Importantly, our approach effectively mitigates reward hacking by explicitly promoting exploration and penalizing trivial, unmodified rewrites. These findings demonstrate that strategy-guided RL not only enhances effectiveness and efficiency but also leads to more transparent and controllable retrieval systems.
6 Limitation
6.1 Evaluation Datasets
One of the key innovations introduced in this paper is the query rewrite strategy or policy. This approach relies on manual analysis to craft rewriting strategies tailored to specific datasets. However, this methodology poses certain limitations. In particular, it is less applicable to domains such as sparse retrieval and SQL-based retrieval, where the query structure is already well-defined and leaves little room for rewriting strategies. Consequently, the benefits of our approach are constrained in these scenarios. Our work is therefore primarily focused on the dense retrieval setting, where more flexibility in query reformulation exists.
@@ A.1 Strategy-Guided Prompting
As referenced in Section 2.2, we present detailed strategies and the complete prompt used to guide our policy model in Table 4 and Figure 5.
@@ A.2 SGP Performance with Contriever
To confirm that our findings are robust and not specific to a single retriever, we replicated the experiment using Contriever Izacard et al. (2021). As detailed in Fig 6, our strategy again demonstrates a significant performance margin over the methods proposed by DMQR-RAGLi et al. (2024).
Figure 6: Performance scaling laws for query rewriting using the Contriever Izacard et al. (2021) retriever across four benchmark datasets: HotpotQA Yang et al. (2018), FEVER Thorne et al. (2018), NFCorpus Boteva et al. (2016), and SciFact Wadden et al. (2020). Each subplot compares the upper-bound performance (NDCG@10) of our SGP against the strategies from Li et al. (2024), evaluated via a Best-of-N methodology across the Qwen3 model series Yang et al. (2025a). The results demonstrate the superior scaling potential and higher performance ceiling of our approach.
A.3 Training Hyperparameters
For all experiments, we fine-tune the Qwen3-4B Yang et al. (2025a) model using Verl Sheng et al. (2024) and GROP Shao et al. (2024), a reinforcement learning framework. We experiment with rollout numbers of 16 and 32. All training is conducted on 8 NVIDIA A100 GPUs. Unless otherwise specified, we use the default hyperparameters provided by Verl.
A.4 Detailed Training and Response Length Curves
@@
A.5 Related Work
The Evolution of Query Rewriting
Reinforcement Learning for Query Rewriting
To alleviate reliance on explicit supervision, recent approaches leverage RL to guide the generative capabilities of LLMs using weaker signals Schulman et al. (2017); Shao et al. (2024); Jiang et al. (2025); Li et al. (2024); Wang et al. (2025b, d, c). However, the application of these powerful algorithms often reveals significant limitations. A common approach is to treat the LLM as a black box, applying on-policy algorithms like PPO without modification Jiang et al. (2025), which overlooks the need for specialized guidance. Other lines of work focus on prompt engineering, but strategies tailored for sparse retrieval in general web search Li et al. (2024) exhibit limited effectiveness when transferred to dense retrieval tasks. Furthermore, even approaches that integrate reward models, such as Wang et al. (2025d), can be flawed; they often rely on an arbitrary, static fusion of scores from multiple fine-tuned reward models, lacking dynamic feedback from the actual retrieval environment. This highlights a fundamental gap: existing RL-based methods frequently neglect the nuanced, domain-specific guidance required to fully exploit the potential of dense retrieval systems. In contrast, our SAGE framework addresses this gap by introducing a novel layer of explicit, human-designed strategic guidance. It achieves this through two core innovations: first, by equipping the agent with a set of fine-grained strategies for query augmentation, and second, by proposing novel reward calculation schemes like SCS and CRS to provide a more nuanced and effective learning signal.
Challenge in RL: Reward Hacking